# Fast data inversion for high-dimensional dynamical systems from noisy measurements

### v2

@@ 2.3 Estimation of the number of latent processes
Estimating a suitable number of factors is crucial to distinguish signals from noise. Various approaches have been developed based on, for instance, information criteria (Bai and Ng,, 2002), the cumulative percentage of variance explained by the factors (Hotelling,, 1933), and the ratio of neighboring eigenvalues of the empirical autocovariance matrix (Lam et al.,, 2011). In our scientific application, the goal is not to select the number of factors, but to provide a suitable model for linking the complex slip propagation process to ground deformation. To enable the variability in the signal to be properly explained by our model, we select the number of latent factors, $d$, by minimizing the difference between the estimated and measured noise variance:
$\hat{d}=\operatorname{argmin}_{d}|({\sigma}^{\text{MMLE}}_{0}(d))^{2}-\sigma_{0}^{2}|,$ (22)
@@
When the variance of the noise is unknown, we use an information criterion to select the number of factors (Bai and Ng,, 2002), which often has the form below:
$\hat{d}=\operatorname{argmin}_{d}\left\{\log(||\mathbf{Y}-\tilde{\mathbf{U}}_{1:d}\tilde{\mathbf{U}}_{1:d}^{T}\mathbf{Y}||_{F})+C(k,n)\right\},$ (23)
@@ 3 Improved computational scalability and efficiency
We discuss the connection to several other approaches, including the dynamic mode decomposition (DMD) (Schmid,, 2010) in Section 3.1, and the generalized probabilistic principal component analysis (GPPCA) (Gu and Shen,, 2020) in Section 3.2. For estimating slips from geodetic data, the network inversion filter (NIF) (Segall and Matthews,, 1997) and its modification (Bartlow et al.,, 2011) are a class of vector autoregressive models, summarized in Section S4 in the supplementary materials. The computational order of different approaches is provided in Table 1.
Methods computational order a noise model closed-form expressions of $\mathbf{U}_{0}$
@@
Table 1: Comparison of different approaches by the computational order, whether a noise model is included, and whether a close-formed expression in estimating the factor loading matrix is available when the factor processes have distinct correlation parameters. For FMOU, we assume $M$ iterations in EM, where $M\approx 10$s often achieve good performance. For DMD, we only consider the cost of obtaining the first $d$ eigenvectors, whereas obtaining the transition coefficient matrix requires extra $O(k^{2}d)$ operations. When each factor contains distinct correlation parameters, GPPCA requires $M_{1}$ iterations and each requires $M_{2}$ steps to optimize the factor loading matrix in the Stiefel manifold. To estimate the slip, an additional SVD operation of a $k\times k^{\prime}$ matrix $\mathbf{G}$ is needed and it only needs to be done once. Thus, the order of FMOU is $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))+\mathcal{O}(Mknd)$ and the order of NIF is $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))+\mathcal{O}(\tilde{M}k^{3}n)$ with $\tilde{M}$ being the number of iterations in NIF for slip estimation.
@@
The DMD is a popular approach in computational mathematics to obtain the reduced-rank representation of nonlinear dynamical systems (Schmid,, 2010; Tu et al.,, 2014). In DMD, one approximates a real-valued output of $k$ dimensions at time $t$ through $\mathbf{y}(t)\approx\mathbf{A}\mathbf{y}(t-1)$, where $\mathbf{A}$ is a $k\times k$ linear transition matrix, estimated by
$\mathbf{\hat{A}}=\operatorname{argmin}_{\mathbf{A}}\lVert\mathbf{Y}_{2:n}-\mathbf{A}\mathbf{Y}_{1:n-1}\rVert=\mathbf{Y}_{2:n}\mathbf{Y}_{1:n-1}^{+},$ (24)
@@
Estimating the leading eigenpairs of $\mathbf{A}$ is crucial for finding the low-rank representation of the system. In Tu et al., (2014), the authors proposed an efficient algorithm named exact DMD, which identifies the nonzero eigenvalues and the corresponding eigenvectors of $\mathbf{A}$ with $\min(\mathcal{O}(kn^{2}),\mathcal{O}(nk^{2}))$ operations. The exact DMD algorithm is summarized in Section S4.1 in the supplementary materials.
A probabilistic model of the DMD was introduced in Gu et al., 2024b , which shows the estimation of $\mathbf{A}$ in DMD is equivalent to the maximum likelihood estimator of $\mathbf{A}$ in the vector autoregressive model $\mathbf{y}(t)=\mathbf{A}\mathbf{y}(t-1)+\bm{\mathbf{\tilde{w}}}(t)$, where $\mathbf{\tilde{w}}(t)\sim\mathcal{MN}(\mathbf{0},\mathbf{\Sigma}_{\mathbf{\tilde{w}}})$ represents the innovation with a positive definition matrix $\mathbf{\Sigma}_{\mathbf{\tilde{w}}}$. However, both DMD and its probabilistic model assume noise-free observation, which is restrictive in practice.
@@
$\mathbf{y}(t)$ $=\mathbf{\bar{z}}(t)+\bm{\epsilon}(t),$ (25)
$\mathbf{\bar{z}}(t)$ $=\mathbf{A}\mathbf{\bar{z}}(t-1)+\bar{\mathbf{w}}(t),$ (26)
@@
Thus the FMOU approach provides a fast solution of a noise-inclusive, probabilistic DMD model with a symmetric transition matrix. This extension drastically improves the DMD when the data contain noises, as will be numerically demonstrated in Section 4. The computational order of DMD and FMOU is summarized in Table 1, and both are fast in practice. When both $k$ and $n$ are large, the FMOU can be faster than the DMD.
@@
The FMOU model is connected to the GPPCA approach (Gu and Shen,, 2020), which assumes each latent factor follows a Gaussian process. However, when the variance and range parameters of the latent processes are distinct, in each iteration of the parameter estimation, one needs to apply an iterative algorithm to estimate the factor loading matrix on the Stiefel manifold (Wen and Yin,, 2013). In total, one needs $M_{1}M_{2}$ iterations, where $M_{1}$ are for numerical optimization of the parameters, each requiring $M_{2}$ iterations to optimize the factor loading matrix. This process can be prohibitively expensive. Furthermore, numerical optimization of $d$ sets of range and variance parameters can also be unstable. In FMOU, we only need $M$ EM iterations, as the estimation of parameters have closed-form expressions. We use Experiment 1 to numerically compare the stability and computational speed of FMOU and GPPCA.
@@
The data are generated by Equations (3)-(4). The orthogonal matrix $\mathbf{U}_{0}\in\mathbb{R}^{k\times d}$ is sampled from the Stiefel manifold with two configurations with $(k=20,d=5)$ and $(k=40,d=10)$, where $d$ is assumed to be known. The latent factors are generated with inputs $t=1,\ldots,n$ for three scenarios with $n\in\{100,200,400\}$. The parameters $\rho_{l}$ and $\sigma_{l}^{2}$ are sampled from $\mbox{Unif}(0.95,1)$ and $\mbox{Unif}(0.5,1)$ for $l=1,...,d$, respectively, and the variance of the noise is $\sigma_{0}^{2}=0.2$. We repeat the simulation $N=20$ times under each configuration.
@@
$\phi_{l}$ $=\text{arccos}\left(\underset{\mathbf{u}\in\mathcal{M}(\mathbf{U}_{0}),\hat{\mathbf{u}}\in\mathcal{M}(\hat{\mathbf{U}}_{0})}{\max}|\mathbf{u}^{T}\hat{\bf u}|\right)=\text{arccos}(|\mathbf{u}_{l}^{T}\hat{\mathbf{u}}_{l}|),$ (27)
such that $\lVert\mathbf{u}\rVert=\lVert\hat{\mathbf{u}}\rVert=1,\mathbf{u}^{T}\mathbf{u}_{l}=\hat{\mathbf{u}}^{T}\hat{\mathbf{u}}_{l}=0$, with ${\bf u}_{l}$ and $\hat{\bf u}_{l}$ being the $l$th column of $\mathbf{U}_{0}$ and $\hat{\mathbf{U}}_{0}$ for $l=1,\ldots,d-1$.
@@
Hence, FMOU achieves tremendous improvements in terms of scalability and efficiency for the model in (3)-(4), compared to GPPCA. In Section 4, we use extensive examples with correctly specified or misspecified scenarios to compare FMOU with other approaches.
@@ 4 Simulated experiments
In this section, we numerically compare our method with alternative approaches by simulation. The experiments are split into two subsections. The scenarios with an estimated factor loading matrix are considered in Section 4.1, whereas Section 4.2 focuses on cases with a fixed factor loading matrix. Both correctly specified models and misspecified models are considered. In Section 4.1, we compared with DMD and two data-driven latent factor models, denoted as LY1 and LY5, introduced in Lam and Yao, (2012). In the LY1 and LY5 methods, the number of latent factors $\hat{d}$ is first estimated by minimizing the ratio of neighboring eigenvalues of $\mathbf{C}:=\sum_{p=1}^{p_{0}}\bm{\hat{\Sigma}}_{y}(p)\bm{\hat{\Sigma}}^{T}_{y}(p)$, with $p_{0}=1$ for LY1 and $p_{0}=5$ for LY5, where $\bm{\hat{\Sigma}}_{y}(p)$ denotes the sample covariance matrix with a lag time $p$. In both methods, the factor loading matrix $\mathbf{U}_{0}$ is estimated by the first ${d}$ eigenvectors of $\mathbf{C}$ corresponding to the largest ${d}$ eigenvalues. Results with an estimated latent factor loading matrix $\hat{d}$ are shown in Section S5 in supplementary materials. In Section 4.2, we construct two additional experiments where the factor loadings are either sampled from the Stiefel manifold or derived from the singular vectors of the Green’s function used in real data, mimicking the slip propagation process. We compare FMOU with NIF (Segall and Matthews,, 1997) for these two experiments. Other than estimation error and computational time, we also record the proportion of the signals and slips covered in $95\%$ posterior credible intervals and the average length of the intervals of the FMOU approach in Tables S2-S6 in supplementary materials, yet these uncertainty measures are not available for other methods, such as DMD.
@@ 4.1 Simulated experiments with estimated factor loadings
(1) Linear diffusion (Carslaw,, 1906). The signal is governed by the partial differential equation $\frac{\partial u(x,t)}{\partial t}=D\frac{\partial^{2}u(x,t)}{\partial x^{2}}$, where $u(x,t)$ represents the concentration of the diffusing material at location $x$ and time $t$, and $D$ is the diffusion coefficient. We follow Lu and Tartakovsky, (2020) to let $D=1$, discretize the spatial domain $[0,1]$ into 500 equally spaced grid points, $u(x,0)=0$, and let a boundary condition be applied at one end with a constant external concentration of 1. The signal is generated over a time interval $t\in[0,0.2]$ using $n=500$ with a numerical solver (Soetaert et al.,, 2010). Three noise variances $\sigma^{2}_{0}\in\{0.01^{2},0.05^{2},0.3^{2}\}$ are tested. (2) Lorenz 96 system (Lorenz,, 1996). The signal follows $\frac{df_{m}(t)}{dt}=(f_{m+1}(t)-f_{m-2}(t))f_{m-1}(t)-f_{m}(t)+F$, for $m=1,\dots,40$ and $F=8$. We generate the signal using the fourth-order Runge Kutta method with a step size $0.01$ for $n=500$ time steps and three noise variances $\sigma_{0}^{2}\in\{0.1,0.5,1\}$ are included to generate observations. We repeat the simulation $N=20$ times for each scenario.
@@
Table 3 presents the average $\text{RMSE}_{m}$ for signal estimation over $N=20$ simulations across various approaches. The results show that FMOU consistently achieves better accuracy under all noise levels, particularly excelling in large-noise scenarios. Figures 4 and 5 provide the comparison between FMOU and the alternative approaches in estimating the signal from noisy observations generated by the linear diffusion equation and Lorenz 96 system, respectively. The signal estimated by FMOU closely aligns with the ground truth, while DMD exhibits larger deviations. Additionally, the 95% posterior credible intervals in Figure 5 correctly cover the true signal for most of the time steps.
Linear diffusion Lorenz 96
$\sigma_{0}^{2}=0.01^{2}$ $\sigma_{0}^{2}=0.05^{2}$ $\sigma_{0}^{2}=0.3^{2}$ $\sigma_{0}^{2}=0.1$ $\sigma_{0}^{2}=0.5$ $\sigma_{0}^{2}=1$
FMOU $\bf{1.67\times 10^{-3}}$ $\bf{6.96\times 10^{-3}}$ $\bf{3.39\times 10^{-2}}$ $\bf{0.185}$ $\bf{0.325}$ $\bf{0.418}$
DMD $9.41\times 10^{-3}$ $4.68\times 10^{-2}$ $2.85\times 10^{-1}$ 0.2800.280{0.280}0.280 0.5710.571{0.571}0.571 $0.770$
LY1 $1.10\times 10^{-1}$ $7.12\times 10^{-2}$ $3.00\times 10^{-1}$ $0.666$ $0.942$ $1.114$
LY5 $1.00\times 10^{-2}$ $5.01\times 10^{-2}$ $3.00\times 10^{-1}$ $1.696$ $2.254$ $2.417$
@@
Table 3: Average of $\text{RMSE}_{m}$ over $N=20$ repeats of Experiment 3.
Figure 4: Noisy observations and predictive signals of the Linear diffusion example in Experiment 3. (a) Observations with large noise ($\sigma_{0}^{2}=0.3^{2}$). (b) Residuals between predictive signals by FMOU and the ground truth, i.e., $u(x,t)-\hat{u}_{\text{FMOU}}(x,t)$. (c) Residuals between predictive signals by DMD and the ground truth, i.e., $u(x,t)-\hat{u}_{\text{DMD}}(x,t)$. Subplots (b) and (c) share the same color key.
Figure 5: Predictive signals by FMOU (dashed blue curves), DMD (dot-dash purple curves), LY1 (long dash orange curves) and LY5 (solid green curves), from one simulation of Lorenz 96 system in Experiment 3 with $\sigma_{0}^{2}=1$. The black circles and solid black curves represent the observations and signals, respectively. The 95% posterior credible intervals by FMOU are graphed as the blue-shaded area. The left and right panels show the 1st and 30th states.

### v3

@@ 2.4 Estimation of the number of latent processes
Estimating a suitable number of factors is crucial to distinguish signals from noise. Various approaches have been developed based on, for instance, information criteria [3], the cumulative percentage of variance explained by the factors [20], and the ratio of neighboring eigenvalues of the empirical autocovariance matrix [23]. In our scientific application, the goal is not to select the number of factors, but to provide a suitable model for linking the complex slip propagation process to ground deformation. To enable the variability in the signal to be properly explained by our model, we select the number of latent factors, $d$, by minimizing the difference between the estimated and measured noise variance:
(22) $\hat{d}=\operatorname{argmin}_{d}|({\sigma}^{\text{MMLE}}_{0}(d))^{2}-\sigma_{0}^{2}|,$
@@
When the variance of the noise is unknown, we follow [3] to use an information criterion to select the number of factors, which often has the form below:
(23) $\hat{d}=\operatorname{argmin}_{d}\left\{\log(||\mathbf{Y}-\tilde{\mathbf{U}}_{1:d}\tilde{\mathbf{U}}_{1:d}^{T}\mathbf{Y}||_{F})+C(k,n)\right\},$
@@ 3 Improved computational scalability and efficiency
We discuss the connection to other approaches, including the dynamic mode decomposition [32], the conventional EM algorithm for vector autoregressive models [34], and the generalized probabilistic principal component analysis [18] in Sections 3.1-3.3, respectively. We also compare the network inversion filter (NIF) [33] and its modification [4] for slip estimation in Section SM4 in supplementary materials. The computational order of different approaches is provided in Table 1.
Methods computational order noise models explicit estimators of $\mathbf{U}_{0}$
@@
Table 1: Comparison of different approaches by the computational order, whether a noise model is included, and whether a closed-form expression in estimating the factor loading matrix is available when the factor processes have distinct correlation parameters. For FMOU, we assume $M$ iterations in EM, where $M\approx 10$s often achieve good performance. For EM-VAR(1), $M^{*}$ is the number of E-M iterations, and the order is when $d=k$, which is the default setting in [36]. For DMD, we only consider the cost of obtaining the first $d$ eigenvectors, whereas obtaining the transition coefficient matrix requires extra $O(k^{2}d)$ operations. When each factor contains distinct correlation parameters, GPPCA requires $M_{1}$ iterations and each requires $M_{2}$ steps to optimize the factor loading matrix in the Stiefel manifold. To estimate the slip, an additional SVD operation of a $k\times k^{\prime}$ matrix $\mathbf{G}$ is needed, and it only needs to be done once. Thus, the order of FMOU is $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))+\mathcal{O}(Mknd)$ and the order of NIF is $\mathcal{O}(min({k}^{2}k^{\prime},{k}(k^{\prime})^{2}))+\mathcal{O}(\tilde{M}k^{3}n)$ with $\tilde{M}$ being the number of iterations in NIF for slip estimation.
@@
The dynamic mode decomposition (DMD) is a popular approach in computational mathematics to obtain the reduced-rank representation of nonlinear dynamical systems [32, 39]. In DMD, one approximates a real-valued output of $k$ dimensions at time $t$ through $\mathbf{y}(t)\approx\mathbf{A}\mathbf{y}(t-1)$, where $\mathbf{A}$ is a $k\times k$ linear transition matrix, estimated by
(24) $\mathbf{\hat{A}}=\operatorname{argmin}_{\mathbf{A}}\lVert\mathbf{Y}_{2:n}-\mathbf{A}\mathbf{Y}_{1:n-1}\rVert=\mathbf{Y}_{2:n}\mathbf{Y}_{1:n-1}^{+},$
@@
Estimating the leading eigenpairs of $\mathbf{A}$ is crucial for low-rank approximation of the system. In [39], the authors proposed the exact DMD algorithm, which computes the nonzero eigenvalues and the corresponding eigenvectors of $\mathbf{A}$ with $\min(\mathcal{O}(kn^{2}),\mathcal{O}(nk^{2}))$ operations. The exact DMD algorithm is summarized in Section SM4.1 in the supplementary materials.
A probabilistic model of the DMD was introduced in [17], which shows the estimation of $\mathbf{A}$ in DMD is equivalent to the maximum likelihood estimator of $\mathbf{A}$ in the vector autoregressive model $\mathbf{y}(t)=\mathbf{A}\mathbf{y}(t-1)+\bm{\mathbf{\tilde{w}}}(t)$, where $\mathbf{\tilde{w}}(t)\sim\mathcal{MN}(\mathbf{0},\mathbf{\Sigma}_{\mathbf{\tilde{w}}})$ represents the innovation with a positive definition matrix $\mathbf{\Sigma}_{\mathbf{\tilde{w}}}$. However, both DMD and its probabilistic model assume noise-free observation, which is restrictive in practice.
@@
(25) $\mathbf{y}(t)$ $=\mathbf{\bar{z}}(t)+\bm{\epsilon}(t),$
(26) $\mathbf{\bar{z}}(t)$ $=\mathbf{A}\mathbf{\bar{z}}(t-1)+\bar{\mathbf{w}}(t),$
@@
Thus, the FMOU approach provides a fast solution of a noise-inclusive, probabilistic DMD model with a symmetric transition matrix. This extension drastically improves the DMD when the data contain noise, as will be numerically demonstrated in Section 4. The computational order of DMD and FMOU is summarized in Table 1, and both are fast in practice. When both $k$ and $n$ are large, the FMOU can be faster than the DMD.
3.2 The FMOU approach accelerates the EM algorithm for vector autoregressive models and enables dimension reduction
The EM algorithm was developed for vector autoregressive models [34]:
(27) $\mathbf{y}(t)$ $=\mathbf{G}_{t}\mathbf{z}(t)+\bm{\epsilon}(t),$
(28) $\mathbf{z}(t)$ $=\bm{\Phi}\mathbf{z}(t-1)+\mathbf{w}(t),$
where $\mathbf{G}_{t}$ is a $k\times d$ known factor loading matrix often specifed as identity matrix, $\bm{\epsilon}(t)\sim\mathcal{MN}(\mathbf{0},\bm{\Sigma}_{\epsilon})$ with covariance $\bm{\Sigma}_{\epsilon}$, $\bm{\Phi}$ is an $d\times d$ coefficient matrix and $\mathbf{w}(t)\sim\mathcal{MN}(\mathbf{0},\bm{\Sigma}_{w})$. Typically, one assumes $\mathbf{G}_{t}=\mathbf{I}_{k}$ with $k=d$ [36]. The EM algorithm was derived for estimating parameters $\{\bm{\Sigma}_{\epsilon},\bm{\Phi},\bm{\Sigma}_{w}\}$ along with the KF. We call this approach the EM algorithm of the vector autoregressive model of order 1 (EM-VAR(1)).
Our FMOU approach has three advantages over this model. First, we developed fast algorithms to estimate the factor loading matrix and enable selecting a reduced dimension $d$, while an identity factor loading matrix is often assumed with $d=k$ in the EM-VAR(1) approach, when $\mathbf{G}_{t}$ is unknown. Second, the KF and RTS smoothers in our FMOU approach only take $\mathcal{O}(nd)$ operations, which is substantially faster than the EM algorithm as it requires $\mathcal{O}(nk^{3})$ for inverting a $k\times k$ matrix at each time step in KF. Third, each factor process of the FMOU model can be set to be stationary, yet the EM-VAR(1) approach cannot guarantee that the latent process is stationary. The computational complexity of EM-VAR(1) model [34, 36] is given in Table 1. We will numerically compare FMOU and EM-VAR(1) in Experiment 3.
@@
The FMOU model is connected to the GPPCA approach [18], which assumes each latent factor follows a Gaussian process. However, when the variance and range parameters of the latent processes are distinct, in each iteration of the parameter estimation, one needs to apply an iterative algorithm to estimate the factor loading matrix on the Stiefel manifold [42]. In total, one needs $M_{1}M_{2}$ iterations, where $M_{1}$ is the number of iterations in numerical optimization of the parameters, each requiring $M_{2}$ iterations to optimize the factor loading matrix. This process can be prohibitively expensive. Furthermore, numerical optimization of $d$ sets of range and variance parameters can also be unstable. In FMOU, we only need $M$ EM iterations, as the estimation of parameters have closed-form expressions. We use Experiment 1 to numerically compare the stability and computational cost of FMOU and GPPCA.
@@
[…]
@@
(29) $\phi_{l}$ $=\text{arccos}\left(\underset{\mathbf{u}\in\mathcal{M}(\mathbf{U}_{0}),\hat{\mathbf{u}}\in\mathcal{M}(\hat{\mathbf{U}}_{0})}{\max}|\mathbf{u}^{T}\hat{\bf u}|\right)=\text{arccos}(|\mathbf{u}_{l}^{T}\hat{\mathbf{u}}_{l}|),$
such that $\lVert\mathbf{u}\rVert=\lVert\hat{\mathbf{u}}\rVert=1,\mathbf{u}^{T}\mathbf{u}_{l}=\hat{\mathbf{u}}^{T}\hat{\mathbf{u}}_{l}=0$, with ${\bf u}_{l}$ and $\hat{\bf u}_{l}$ being the $l$th column of $\mathbf{U}_{0}$ and $\hat{\mathbf{U}}_{0}$ for $l=1,\ldots,d-1$. A smaller largest principal angle indicates a better estimation.
@@
Hence, FMOU achieves tremendous improvements in terms of scalability and efficiency for the model in (3)-(4), compared to GPPCA. In Sections 4 and 5, we use extensive simulated and real examples with correctly specified or misspecified scenarios to compare FMOU with alternative approaches. The numerical results are computed by macOS Mojave system with an 8-core Intel i9 processor running at 3.60 GHz and 32 GB of RAM.
@@ 4 Simulated experiments
In this section, we numerically compare our method with alternative approaches by simulation. The experiments are split into two subsections. The scenarios with an estimated factor loading matrix are considered in Section 4.1, whereas Section 4.2 focuses on cases with a fixed factor loading matrix. Both correctly specified models and misspecified models are considered. In Section 4.1, we compared with DMD and two data-driven latent factor models, denoted as LY1 and LY5, introduced in [22]. In the LY1 and LY5 methods, the number of latent factors $\hat{d}$ is first estimated by minimizing the ratio of neighboring eigenvalues of $\mathbf{C}:=\sum_{p=1}^{p_{0}}\bm{\hat{\Sigma}}_{y}(p)\bm{\hat{\Sigma}}^{T}_{y}(p)$, with $p_{0}=1$ for LY1 and $p_{0}=5$ for LY5, where $\bm{\hat{\Sigma}}_{y}(p)$ denotes the sample covariance matrix with a lag time $p$. In both methods, the factor loading matrix $\mathbf{U}_{0}$ is estimated by the first ${d}$ eigenvectors of $\mathbf{C}$ corresponding to the largest ${d}$ eigenvalues. Results with an estimated latent factor loading matrix $\hat{d}$ are shown in Section SM5 in supplementary materials. For Experiment 3 in Section 4.1, we also include EM-VAR(1) [34] with the default setting $\mathbf{G}_{t}=\mathbf{I}_{k}$ in Equation (27) for estimating the mean from noisy observations. The non-orthogonal eigenvectors of the $\Phi$ in Equation (28) play a similar role as the factor loading matrix in the FMOU model, which is estimated by the data. We do not include EM-VAR(1) in other simulated examples as they are too expensive to compute. In Section 4.2, we construct two additional experiments where the factor loadings are either sampled from the Stiefel manifold or derived from the singular vectors of the Green’s function used in real data, mimicking the slip propagation process. We compare FMOU with NIF [33] for these two experiments. Other than estimation error and computational time, we also record the proportion of the signals and slips covered in $95\%$ posterior credible intervals and the average length of the intervals of the FMOU approach in Tables SM2-SM6 in supplementary materials, yet these uncertainty measures are not available for other methods, such as DMD.
@@ 4.1 Simulated experiments with estimated factor loadings
(1) Linear diffusion [8]. The signal is governed by the partial differential equation $\frac{\partial u(x,t)}{\partial t}=D\frac{\partial^{2}u(x,t)}{\partial x^{2}}$, where $u(x,t)$ represents the concentration of the diffusing material at location $x$ and time $t$, and $D$ is the diffusion coefficient. We follow [24] to let $D=1$, discretize the spatial domain $[0,1]$ into $300$ equally spaced grid points, $u(x,0)=0$, and let a boundary condition be applied at one end with a constant external concentration of 1. The signal is generated over a time interval $t\in[0,0.2]$ using $n=300$ with a numerical solver [35]. Three noise variances $\sigma_{0}^{2}\in\{0.01^{2},0.05^{2},0.3^{2}\}$ are tested. (2) Branin function [29]. The underlying signal is generated by the Branin function, which takes a two-dimensional input $(x_{1},x_{2})$ and yields a scalar output given by $f(x_{1},x_{2})=a(x_{2}-bx_{1}^{2}+cx_{1}-r)^{2}+s(1-t)\cos(x_{1})+s$, where $x_{1}\in[-5,10]$ and $x_{2}\in[0,15]$. The recommended parameter values are used: $a=1,b=\frac{5.1}{4\pi^{2}},c=\frac{5}{\pi},r=6,s=10$ and $t=\frac{1}{8\pi}$. When generating the signal, the input domain is uniformly discretized into a $300\times 300$ grid. Noisy observations are then obtained via $y(x_{1},x_{2})=f(x_{1},x_{2})+\epsilon$, where $\epsilon$ represents independent Gaussian noise with variance $\sigma^{2}_{0}$. We test the effects of varying noise variances, specifically $\sigma_{0}^{2}\in\{1,5^{2},20^{2}\}$.
@@
Table 3 presents the average $\text{RMSE}_{m}$ and computational time for signal estimation over $N=20$ simulations across various approaches. The results show that FMOU consistently achieves better accuracy under all noise levels, particularly excelling in large-noise scenarios. Moreover, FMOU is much faster over the EM-VAR(1) method with the default setting [36] as the FMOU enables dimension reduction of the latent factors and avoids inverting the covariance matrix of the observations at each time in the Kalman filter. Figure 4 provides the comparison between FMOU and DMD in estimating the signal from noisy observations generated by the linear diffusion equation and Branin function, respectively. The signal estimated by FMOU closely aligns with the ground truth, while DMD exhibits larger deviations. We highlight that the convergence of the log likelihood in the FMOU approach only takes around 5-10 iterations in EM algorithm shown in Figure 4 (d) and (h).
Linear diffusion Branin function
RMSEm RMSEm RMSEm Avg RMSEm RMSEm RMSEm Avg
$\sigma^{2}_{0}=0.01^{2}$ $\sigma^{2}_{0}=0.05^{2}$ $\sigma^{2}_{0}=0.3^{2}$ Time $\sigma^{2}_{0}=1$ $\sigma^{2}_{0}=5^{2}$ $\sigma^{2}_{0}=20^{2}$ Time
FMOU $\bf{0.0024}$ $\bf{0.0083}$ $\bf{0.038}$ $0.56$ $\bf{0.14}$ $\bf{0.60}$ $\bf{2.2}$ $\bf{0.23}$
DMD $0.010$ $0.047$ $0.29$ $\bf{0.44}$ 0.790.79{0.79}0.79 4.54.5{4.5}4.5 $19$ $0.41$
LY1 $0.10$ $0.050$ $0.30$ $3.6$ $2.4$ $11$ $20$ $3.5$
LY5 $0.010$ $0.050$ $0.30$ $18$ $5.9$ $5.8$ $20$ $17$
EM-VAR(1) $0.0098$ $0.050$ $0.30$ $672$ $4.1$ $5.5$ $60$ $807$
@@
Table 3: Average $\text{RMSE}_{m}$ and computational time (in seconds) for Experiment 3. $\text{RMSE}_{m}$ is averaged over $N=20$ repeats per noise level. The computational time is averaged over three noise levels and $N=20$ repeats.
Figure 4: (a) True mean generated by linear diffusion. (b) Predictive mean by FMOU, where observations are generated with $\sigma_{0}^{2}=0.05^{2}$. (c) Predictive mean by DMD (i.e., $\hat{u}_{\text{DMD}}(x,t)$) where observations are generated with $\sigma_{0}^{2}=0.05^{2}$. (d) The convergence of the log likelihood in the EM algorithm of the FMOU approach. (e) True mean generated by the Branin function. (f) Predictive mean by FMOU, where observations are generated with $\sigma_{0}^{2}=25$.(g) Predictive mean by DMD, where observations are generated with $\sigma_{0}^{2}=25$. (h) The convergence of the log likelihood in the EM algorithm of the FMOU approach.